OpenStudy (anonymous):

Saturday morning there were 7 adults and 20 kids at Pump and Jump Land. The morning cash registers show $236 were collected. In the afternoon there were 21 adults and 50 kids for a total amount of $618. What is the cost for 1 adult?

OpenStudy (anonymous):

\[\Large \begin{array}{l} 7x + 20y = 236\\ 21x + 50y = 618 \end{array}\] Solve system.

hero (hero):

7a + 20k = 236 21a + 50k = 618

OpenStudy (anonymous):

i can eliminate the x value right

hero (hero):

You can eliminate whatever you want. As long as you get the right answer

OpenStudy (anonymous):

i got y =-9 and x=59.43

OpenStudy (anonymous):

\[\Large \begin{array}{l} \left[ {\begin{array}{*{20}{c}} 7&{20}\\ {21}&{50} \end{array}} \right] = \left[ {\begin{array}{*{20}{c}} {236}\\ {618} \end{array}} \right]\\ \left[ {\begin{array}{*{20}{c}} 7&{20}\\ 0&{ - 10} \end{array}} \right] = \left[ {\begin{array}{*{20}{c}} {236}\\ { - 90} \end{array}} \right]\begin{array}{*{20}{c}} {{R_1}}\\ {{R_2} - 3 \cdot {R_1}} \end{array}\\ y = \frac{{ - 90}}{{ - 10}} = 9\\ x = \frac{{236 - 9 \cdot 20}}{7} = {\rm{8}} \end{array}\]

OpenStudy (mr.math):

Hey fewscrewsmissing! What's the size of your font? :P

OpenStudy (anonymous):

Large =) I just add \Large to the first line of latex inside the opening brackets.

OpenStudy (mr.math):

I think you have too add "large" at least three times to get a font that huge :P

OpenStudy (anonymous):

Haha it's "Large" with a capital L. large > Large > LARGE I think (case-sensitive).

OpenStudy (anonymous):

Er you know what I mean ;)

OpenStudy (mr.math):

Yeah, I know how to use it ;)